OpenStudy (perl):

the maximum value of 4sinx - 3 cos x , on [ 0 , 2pi] , WITHOUT A CALCULATOR

myininaya (myininaya):

\[f'(x)=4\cos(x)+3\sin(x)=0 \text{ solve for x}\]

OpenStudy (perl):

ok i got tan x = -3/4

OpenStudy (perl):

errr, tan x = -4/3

myininaya (myininaya):

right

OpenStudy (perl):

arctan (-4/3) = x , now i cant use calculator

OpenStudy (perl):

lagrange, did it work ?

OpenStudy (perl):

myin, WHAT ARE THE SOLUTIONS, max and min

myininaya (myininaya):

the exact x=arctan(-4/3)

myininaya (myininaya):

now you need to find out if it is a max or min

myininaya (myininaya):

there is also another x=arctan(-4/3)+180

myininaya (myininaya):

i will show you how to plug in one of these say we look at x=arctan(-4/3)+180 and we want to evaluate our function f at this x we would get f(arctan(-4/3)+180)=4sin(arctan(-4/3)+180) - 3 cos(arctan(-4/3)+180) \[=4[\sin(\arctan(-4/3))\cos(180)+\cos(\arctan(-4/3))\sin(180)] \] \[-3[cos(arctan(-4/3))cos(180)-sin(arctan(-4/3))sin(180)] \] \[=4[\sin(\arctan(-4/3)(-1)+\cos(\arctan(-4/3))(0)]\] \[-3[\cos(\arctan(-4/3)(-1)-\sin(\arctan(-4/3))(0)]\] \[=-4\sin(\arctan(-4/3))+3 \cos(\arctan(-4/3))\] so now we need to evaluate sin(arctan(-4/3)) and cos(arctan(-4/3)) remember if y=arctan(-4/3) => tan(y)=-4/3  we can find the hyp. by using Pythagorean thm and we get 5 for the hyp so sin(y)=-4/5 cos(y)=3/5 so continuing the evaluation we get \[=-4(-4/5)+3(3/5)=\frac{16+9}{5}=\frac{25}{5}=5\]

myininaya (myininaya):

now you should evaluate f(arctan(-4/3))

myininaya (myininaya):

and don't forget to plug in the endpoints

OpenStudy (perl):

ok thanks so much

OpenStudy (anonymous):

wanna fight again?

myininaya (myininaya):

did you understand what i did above in that triangle that angle was suppose to be labeled y

myininaya (myininaya):

oops +pi not 180

OpenStudy (anonymous):

\[[0,2\pi]\]

myininaya (myininaya):

satellite makes everything such a big deal

myininaya (myininaya):

he is saying that they want theta in terms of radians

myininaya (myininaya):

so instead of saying +180 say +pi

OpenStudy (anonymous):

no i am saying they want x in terms of numbers

myininaya (myininaya):

i mean x in terms of radians

OpenStudy (anonymous):

and in any case if you are using sine and cosine as a function of degrees, then the derivative of sine is not ( i repeat not) cosine!

OpenStudy (anonymous):

x is neither a degree nor a radian nor any other specific unit. x is a number as in \[f(x)=x^2\]

myininaya (myininaya):

satellite an angle is measured in degrees or radians why do you say these aren't numbers?

myininaya (myininaya):

30 degrees is not a number?

myininaya (myininaya):

pi/6 is not a number?

OpenStudy (anonymous):

in any case myininays is of course right, just forget the damned degrees unless you are solving a triangle using degrees as your angle measure

OpenStudy (anonymous):

i did not say 30 is not a number. what i am trying to say is that sine and cosine are functions of real numbers, and as such, they only correspond to the functions of angles if the angles are measured in radians.

myininaya (myininaya):

why can't the angles be measured in degrees?

OpenStudy (anonymous):

fine, but then the functions are functions of angles measured in degrees, and do not correspond to the functions as functions of numbers. so in particular the derivative of sine is not cosine if you are thinking of these as functions with domain angles measured in degrees rather than numbers.

myininaya (myininaya):

then what is derivative of sine?

myininaya (myininaya):

if the angle is measured in degrees?

OpenStudy (anonymous):

you can see this clearly from the graph. graph sine and you will see that the x - axis is labeled with numbers like pi, 2pi etc. they are real valued functions with domain all real numbers and period 2 pi

OpenStudy (anonymous):

jesus i worked it out once and it is a pain. you have to use the chain rule.

myininaya (myininaya):

lol ok satellite

myininaya (myininaya):

yes that makes sense

OpenStudy (anonymous):

whew.

OpenStudy (anonymous):

i will be quiet now (until the next time...)

myininaya (myininaya):

i think perl said he gots it from where i left off right?

OpenStudy (anonymous):

actually, not th that you asked, but we can do this problem without calc

myininaya (myininaya):

satellite never likes mine ways

OpenStudy (anonymous):

you can rewrite this as \[5 \sin(x-\tan^{-1}(\frac{3}{4}))\] which obviously has a max of 5.

OpenStudy (anonymous):

just thought i would mentions it. only way i know this is that someone here asked how to write \[a\cos(x)+b\sin(x)\] as a single trig function and looked it up day before yesterday

myininaya (myininaya):

\[\sqrt{a^2+b^2}(\frac{a}{\sqrt{a^2+b^2}}\cos(x)+\frac{b}{\sqrt{a^2+b^2}}\sin(x))\] \[\sqrt{a^2+b^2}(\sin(C+x)), \text{where } C=\tan^{-1}(\frac{a}{b})\]

myininaya (myininaya):

\[4\sin(x)-3\cos(x)=\sqrt{(4)^2+(-3)^2}\sin(\tan^{-1}(\frac{-3}{4})+x)\]

OpenStudy (anonymous):

Attached is a plot from x=0 to x=2 pi + 1

myininaya (myininaya):

remember tan(x) is an add function so tan(-x)=-tan(x) so if tan(-x)=y then -tan(x)=y => arctan(y)=-x , arctan(-y)=x => -arctan(y)=x so you could write the expression like satellite has

OpenStudy (anonymous):

add function lol

myininaya (myininaya):

oops

OpenStudy (anonymous):

just teasing i know what you meant. did you go trick or treating?

myininaya (myininaya):

no you knew what i meant after i said tan(-x)=-tan(x) right? :)

OpenStudy (anonymous):

yes. why not? did kids come to the house?